the sum of two consecutive integers is 119.What are he integers ?
1st integer=59
2nd ineger=60

Answer by checkley77(12844)

X+(X+1)=119
2X+1=119
2X=118-1
2X=118
X=118/2
X=59 FOR THE SMALLER INTEGER.
59+1=60 FOR THE LARGER INTEGER.